A magazine reported that 6% of American drivers read the newspaper while driving. If 300  drivers are selected at random, find the probability that exactly 25 say they read the Newspaper while driving.


1
Expert's answer
2021-05-26T03:45:23-0400

Here

p = 0.06

q = 0.94 and

n = 300


Step1: Check whether a normal approximation can be used

"np=300\times 0.06=18\\nq=300\times 0,94=292"


Since "np\geq5\ \ and \ \ nq\geq5" , we can use the normal distribution.


Step2: To find the mean and standard deviation,

"Mean\ \ \mu=np=18\\Standard\ Deviation\ \ \sigma=\sqrt{npq}=\sqrt{300(0.06)(0.94)}=4.11"


Step3: Write the probability notation. "P(X=25)"


Step4: Rewrite using the continuity correction factor.

"P(24.5<X<25.5)"


Step5: Find the corresponding z values.

"z=\dfrac{24.5-18}{4.11}=1.58\ \ \ ,z=\dfrac{25.5-18}{4.11}=1.82"



Step6: Find the solution.

The area between the two z value is

"0.9656-0.9429 = 0.0227, or \ 2.27 \%"

Hence, the probability that exactly 25 people read the newspaper while driving is 2.27%